OpenStudy (anonymous):

f(x)=x^2-6x-3 vertex? equation of the line symmetry? max/min of f(x) is the value of, f(3)=-12 a min or max? graph?

OpenStudy (anonymous):

I think it is (3,-12) f(3)=-12 -12 minimum

OpenStudy (anonymous):

vertex: (-b/(2a), f(-b/(2a))) line of symmetry: x=-b/(2a) max/min: f(-b/(2a))